OpenStudy (anonymous):

evaluate f(x)=3sin+3cos1/2x when f(x)=x^2-2

OpenStudy (anonymous):

oh i wrote that wrong it is evaluate f(x)=3sin+3cos1/2x at x=pi/3

OpenStudy (anonymous):

sorry :(

zepdrix (zepdrix):

Ah ok now it makes a little more sense :) What is the argument of the sine function?

OpenStudy (anonymous):

what does that mean

zepdrix (zepdrix):

The stuff inside of the function is called the "argument" :) What is suppose to go inside of the sine function, because you have it empty right now. You forgot to write something there.

zepdrix (zepdrix):

3sin(?)

OpenStudy (anonymous):

oh x sorry lol

zepdrix (zepdrix):

So you just plug in pi/3 everywhere you see an x in the problem. What part are you stuck on? Do you remember your special angles? :D

OpenStudy (anonymous):

so is it \[3\sqrt{3}\]

zepdrix (zepdrix):

 Yayyy good job \c:/